OpenStudy (anonymous):

A restaurant has 45 tables. Each table seats 4 people. The manager has 4 employees to prepare the tables. The chart shows how each employee works. Use the information to answer the question Chart2 Employee Time to set 4 places (min) Stacie 5 Jeff 3 Tiffany 2 Nick 4 If everyone starts working at the same time, how quickly can the employees get the tables ready for dinner? The first pair to finish their job should help the other pair to finish the tables. Explain your answ

OpenStudy (anonymous):

There is another chart for folding napkins employees time to fold 4 napkins stacie 3 jeff 4 tiffany 5 nick 3.5

OpenStudy (anonymous):

this is how i interpreted the problem, and what i believe is correct. well since their is 45 tables and each table 4 people; is be 180 places, if we count to see how many places can be done by the min, so if we add 5 + 3 + 2 + 4 = 14. so every min it would be 14, so 180 / 14 = 12. 8571; which would be approximately the time taken. For the napkins it would be the same, 3 + 4 +5 + 3.5 = 15.5 180/15.5 = 11.61 mins.